# Differentially Private Matchings

### v1

@@ 8.1 Privacy Proof
Algorithm 9 is $\varepsilon$-node differentially private.
@@
Finally, for each of the $T$ rounds, we have that Algorithms 9, 9 and 9 can be analyzed as an instance of the Laplace mechanism with privacy parameter $\varepsilon^{\prime}$ and query $c_{0}$. In a given iteration, each bidder makes only a single bid so the sensitivity of the query is $1$, implying that this part is $\varepsilon^{\prime}$-differentially private. Applying the concurrent composition theorem, we have that Algorithms 9, 9 and 9 is $T\varepsilon^{\prime}$-differentially private. Finally, by concurrent composition, the entire algorithm is $T\varepsilon^{\prime}+T\varepsilon^{\prime}+T\varepsilon^{\prime}/\eta\leq\varepsilon$-differentially private. ∎
8.2 Utility Proof
Lemma 8.2.
Proof.
$(s-36c\log(n)/\varepsilon^{\prime})N-Q.$
$N\leq(\text{OPT}+Q)/(s-36c\log(n)/\varepsilon^{\prime}).$
Thus, the total number of bidders who were unmatched is at most
$B\geq T\cdot(Q-36c\log(n)/\varepsilon^{\prime}).$ (1)
Combining our two estimates in Eqs. 1 and 2, we have
implying that
@@ 8.2 Utility Proof
If the supply is at least $s\geq 144c\log(n)/\eta\varepsilon^{\prime}$, Algorithm 9 outputs a $(1+\eta)$-approximate maximum (one-sided) $s$-matching with probability at least $1-1/n^{c}$ for constant $c\geq 1$.
Proof.
$w_{\mu(v),v}-p_{\mu(v)}\geq w_{\mu^{*}(v),v}-p_{\mu^{*}(v)}-\eta.$
Next, observe that if a good $u$ has $p_{u}>0$, this means that at least $s-18c\log(n)/\varepsilon^{\prime}$ bidders were (temporarily) matched with it, due to Algorithm 9. This directly implies that the number of bidders matched with $u$ at the termination is at least $s-36c\log(n)/\varepsilon^{\prime}$, because goods only unmatch with bidders until they are matched with $s-36c\log(n)/\varepsilon^{\prime}$ bidders (Line 9). Thus, there can be at most $\text{OPT}/(s-36c\log(n)/\varepsilon^{\prime})$ goods with $p_{u}>0$. For each of these goods, we have $p^{*}_{u}\cdot N_{u}^{*}-p_{u}\cdot N_{u}\leq s$, so we have
@@ 9.1 Lower Bound for Implicit Matchings
Let $\mathcal{A}$ be an algorithm which satisfies $\varepsilon$-edge DP and outputs an implicit solution which is a $(1+\eta)$-approximate maximal matching with probability at least $1-\beta$, with $\beta\leq 1/(16e^{4\varepsilon})$. Then the degree of this implicit solution is at least $\Omega(\frac{1}{\varepsilon}\log(1/\beta))$ with probability at least $1/2$.
@@
$e^{|T|\varepsilon}2e^{2\varepsilon}\beta$ $=e^{\frac{1}{2\varepsilon}\log(1/\beta)\varepsilon}2e^{2\varepsilon}\beta=(1/\beta)^{1/2}2e^{2\varepsilon}\beta=\beta^{1/2}2e^{2\varepsilon}\leq 1/2.$
Thus with probability at most $1/2$, when we run $\mathcal{A}$ on $G^{\prime}$ the implicit solution we get includes at most $|T|/2$ edges from $r$ to $T$. Thus with probability at least $1/2$, when we run $\mathcal{A}$ on $G^{\prime}$ the implicit solution we get includes at least $|T|/2$ edges from $r$ to $T$. Since all of those edges are also edges of $G^{\prime}$, this means that the degree of the implicit solution is at least $|T|/2=\Theta\left(\frac{1}{\varepsilon}\log(1/\beta)\right)$ with probability at least $1/2$. ∎
@@ 9.2 Lower Bound for Node DP Matching Sparsifiers
Acknowledgments
Felix Zhou acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC).
@@ Acknowledgments
Appendix A Proof of Lemma 2.14
Proof of Lemma 2.14.

### v2

@@ 10.2.1 Bounded Arboricity Sparsifiers for $b$-Matchings
Let $G$ be a graph of arboricity at most $\alpha$ and $\Lambda_{b}\coloneqq 5(1+\nicefrac{{5}}{{\eta}})\cdot 2\alpha+(b-1)$ for some $\eta\in(0,1]$. Suppose $H$ is obtained by marking an arbitrary set of $\min(\deg(v),\Lambda_{b})$ edges incident to each vertex $v$ and taking the subgraph consisting of all edges marked by both endpoints. Then if $\mu_{b}(\cdot)$ denotes the size of a maximum $b$-matching of the input graph,
$\mu_{b}(H)\leq\mu_{b}(G)\leq(1+\eta)\mu_{b}(H).$
In particular, any $(\beta,\zeta)$-approximate maximum $b$-matching for $H$ is an $(\beta(1+\eta),\zeta(1+\eta))$-approximate matching of $G$. Moreover, this holds for $H=\textsc{ContractionSparsify}_{\pi}(G,\Lambda_{b})$ where $\pi\in P_{\binom{n}{2}}$ is a total order over unordered vertex pairs.
[…]
@@
We argue by induction on $b$. The base case of $b=1$ is guaranteed by Theorem 6.2. Suppose inductively the statement holds for $b-1$ for some $b\geq 2$ so that our goal is to extend it to the case $b$.
Let $M_{1}\cup\dots M_{b}$ be any $b$-matching of $G$ decomposed into $b$ disjoint matchings. Let $H$ denote a candidate sparsifier obtained by marking an arbitrary set $T(v)$ of at least $\min(\deg_{G}(v),\Lambda_{b})$ edges incident to each vertex $v$, and taking the subgraph $H=(V,E_{H})$ consisting of all edges marked by both endpoints, i.e. $E_{H}\coloneqq\{uv\in E(G):u\in T(v),v\in T(u)\}$.
Consider $G_{1}\coloneqq G-(M_{2}\cup\dots\cup M_{b})$ and note that $M_{1}$ is a matching in $G_{1}$. We claim that $E_{H}\cap E(G_{1})$ contains a matching of size at least $\frac{1}{1+\eta}|M_{1}|$. Indeed, let
$\forall v\in V,T_{1}(v)\coloneqq\{u\in T(v):uv\notin M_{2}\cup\dots\cup M_{b}\},\qquad E_{1}\coloneqq\{uv\in E:u\in T_{1}(v),v\in T_{1}(u)\}\,.$
Note that each $T_{1}(v)$ contains at least $\min(\deg_{G_{1}}(v),\Lambda_{1})$ neighbors of $v$ within $G_{1}$. Furthermore, we have $E_{1}\subseteq E_{H}$. Thus by the base case, $E_{1}$ and hence $H$ contains a $(1+\eta)$-approximate maximum matching $M_{1}^{\prime}$ of $G_{1}$.
Remark that $M_{1}^{\prime}\cup M_{2}\cup\dots\cup M_{b}$ is a $b$-matching within $G$ by construction. Consider $G_{-1}\coloneqq G-M_{1}^{\prime}$, so that $M_{2}\cup\dots\cup M_{b}$ is a $(b-1)$-matching within $G_{-1}$. We claim that $E_{H}\cap E(G_{-1})$ contains a $(b-1)$-matching of size at least $\frac{1}{1+\eta}|M_{2}\cup\dots\cup M_{b}|$. To prove this claim, we will rely on the induction hypothesis. Define
$\forall v\in V,T_{-1}(v)\coloneqq\{u\in T(v):uv\notin M_{1}^{\prime}\},\qquad E_{-1}\coloneqq\{uv\in E:u\in T_{-1}(v),v\in T_{-1}(u)\}\,.$
Then each $T_{-1}(v)$ contains at least $\min(d_{G_{-1}}(v),\Lambda_{b-1})$ neighbors of $v$ in $G_{-1}$. Furthermore, $E_{-1}\subseteq(E_{H}\setminus M_{1}^{\prime})$ by construction. Thus by the induction hypothesis, $E_{-1}$ and hence $H$ contains a $(1+\eta)$-approximate maximum $(b-1)$-matching $N$ of $G_{-1}$.
We have identified disjoint edge subsets $M_{1}^{\prime},N\subseteq E_{H}$ such that $M_{1}^{\prime}$ is a matching in $G$ with size at least $\frac{1}{1+\eta}|M_{1}|$ and $N$ is a $(b-1)$-matching in $G-M_{1}^{\prime}$ with size at least $\frac{1}{1+\eta}|M_{2}\cup\dots\cup M_{b}|$. Thus $M_{1}^{\prime}\cup N$ is a $b$-matching in $G$ with size at least
$|M_{1}^{\prime}\cup N|=|M_{1}^{\prime}|+|N|\geq\frac{1}{1+\eta}|M_{1}\cup\dots\cup M_{b}|\,.$
By the arbitrary choice of $M_{1}\cup\dots\cup M_{b}$, we conclude the proof. ∎
10.3 Arbitrary Edge-Order Continual Release $b$-Matchings
Similarly, we can extend our $b$-matching algorithms to the continual release setting. Here we state the guarantees for an $\varepsilon$-edge DP algorithm for arbitrary edge-order stream, but our algorithms also extend to the node-DP setting.
@@ 10.3 Arbitrary Edge-Order Continual Release $b$-Matchings
For $\eta\in(0,1)$ and $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\Omega\left(\frac{\log^{2}(n)}{\eta^{3}\varepsilon}\right),$ there is an $\varepsilon$-edge DP algorithm that outputs a sequence of implicit $b$-matchings in the billboard model such that each $b$-matching is a $(2+\eta)$-approximation with respect to the maximum $b^{\prime}$-matching size, with probability at least $1-\frac{1}{n^{c}}$.
@@ 11.1 Lower Bound for Implicit Matchings
Let $\gamma>0$ be arbitrary. Let $\mathcal{A}$ be an algorithm that satisfies $\varepsilon$-edge DP and outputs an implicit solution that has the size of a $\gamma$-approximate maximal matching with probability at least $1-\beta$, with $\beta\leq 1/(16e^{4\varepsilon})$. Then the degree of this implicit solution is at least $\Omega(\frac{1}{\varepsilon}\log(1/\beta))$ with probability at least $1/2$, regardless of the value of $\gamma$.
@@
$e^{|T|\varepsilon}2e^{2\varepsilon}\beta$ $=e^{\frac{1}{2\varepsilon}\log(\nicefrac{{1}}{{\beta}})\varepsilon}2e^{2\varepsilon}\beta=\beta^{-\frac{1}{2}}2e^{2\varepsilon}\beta=\beta^{\frac{1}{2}}2e^{2\varepsilon}\leq\frac{1}{2}.$
Thus with probability at most $\nicefrac{{1}}{{2}}$, when we run $\mathcal{A}$ on $G^{\prime}$ the implicit solution we get includes at most $\nicefrac{{|T|}}{{2}}$ edges from $r$ to $T$. Thus with probability at least $\nicefrac{{1}}{{2}}$, when we run $\mathcal{A}$ on $G^{\prime}$ the implicit solution we get includes at least $\nicefrac{{|T|}}{{2}}$ edges from $r$ to $T$. Since all of those edges are also edges of $G^{\prime}$, this means that the degree of the implicit solution is at least $\nicefrac{{|T|}}{{2}}=\Theta\left(\frac{1}{\varepsilon}\log(\nicefrac{{1}}{{\beta}})\right)$ with probability at least $\nicefrac{{1}}{{2}}$. ∎
@@ 11.2 Lower Bound for Node DP Matching Sparsifiers
Appendix A Deferred Preliminaries
A.1 Concentration Inequalities
Lemma A.1.
Theorem A.2 (Multiplicative Chernoff Bound).
  1. 1.
  2. 2.
A.2 Proof of Lemma 2.14
We now restate and prove Lemma 2.14. See 2.14
Proof.
@@ A.1 Concentration Inequalities
A.3 Multi-Response Sparse Vector Technique
[…]
Theorem A.3 (Theorem 2 in [lyu2017understanding]).
Algorithm 10 is $\varepsilon$-DP.
2 Class SVT($\varepsilon,\Delta,c$)
3       $\varepsilon_{1},\varepsilon_{2}\leftarrow\nicefrac{{\varepsilon}}{{2}}$
4       $\rho\leftarrow\text{Lap}(\nicefrac{{\Delta}}{{\varepsilon_{1}}})$
5       $\operatorname{count}\leftarrow 0$
6
7       Function ProcessQuery($f_{t}(D),\tau_{t}$)
8             if $\operatorname{count}>c$ then
9                   return “abort”
10
12                   return “above”
13                   $\operatorname{count}\leftarrow\operatorname{count}+1$
14
15            else
16                   return “below”
17
18
19
Algorithm 10 Sparse Vector Technique
Appendix B Lower Bound for Explicit Fractional Matchings
We now prove Theorem 1.2 which is restated below for convenience: in the basic 0/1 weight $(\varepsilon,\delta)$-edge DP model (Definition 2.7), the required error is essentially linear, even if we allow for the algorithm to output a fractional $b$-matching for very large $b$. This proof is essentially the same as the proof of Theorem 1.1, but we give it explicitly for completeness.
See 1.2
Given a graph $G=(V,E)$, let $\nu^{f}(G)$ denote the size of the largest fractional matching in $G$ Given a (randomized) algorithm $\mathcal{A}$ which outputs some fractional $b$-matching $X_{G}^{\mathcal{A}}$ when run on input graph $G=(V,E)$, i.e. $X_{G}^{\mathcal{A}}\in[0,1]^{\binom{V}{2}}$ is a random vector satisfying $\sum_{v\neq u}X_{G,uv}^{\mathcal{A}}\leq b$ for all $u\in V$ with probability 1, let $\nu_{\mathcal{A}}^{f}(G)=\nu^{f}((V,X_{G}^{\mathcal{A}}\big{|}_{E}))$ be the size of a maximum fractional matching in the graph restricted to edges that are in $G$.
We now begin our proof of Theorem 1.2. At a very high level, we will give a collection of input graphs and argue that any differentially private algorithm must be a poor approximation on at least one of them. This is in some ways similar to the standard “packing arguments” used to prove lower bounds for differentially private algorithms (see, e.g., [DMN23]), but instead of using a packing / averaging argument we will instead use a symmetry argument. We will argue that without loss of generality, any DP algorithm for our class of inputs must be highly symmetric in that $\text{E}[X_{G,uv}^{\mathcal{A}}]$ for each non-edge $uv\notin E$ is identical. This makes it easy to argue that high-quality algorithms cannot satisfy differential privacy.
We begin with a description of our inputs and some more notation. Given $n$ nodes $V$ for an even integer $n>0$, let $\mathcal{M}$ denote the set of all perfect matchings on $V$. Let $\mathcal{A}$ be an $(\varepsilon,\delta)$-differentially private algorithm which always returns a fractional $b$-matching $X_{G}^{\mathcal{A}}$ on $V$ as defined above. For $G\in\mathcal{M}$ and $uv\in\binom{V}{2}$, let $\text{E}[X_{G,uv}^{\mathcal{A}}]$ denote the expected fractional value that $\mathcal{A}(G)$ assigns to $uv$. Let $\nu_{\mathcal{A}}^{f}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}}^{f}(G)]$ denote the worst-case expected utility of $\mathcal{A}$ over $\mathcal{M}$. Note that for $G\in\mathcal{M}$, every subset of $E(G)$ is a matching. By linearity of expectations, this is equivalent to saying that $\sum_{uv\in E(G)}{\text{E}[X_{G,uv}^{\mathcal{A}}]}\geq\nu_{\mathcal{A}}^{f}$ for all $G\in\mathcal{M}$.
Lemma B.1.
If $\mathcal{A}$ satisfies $(\varepsilon,\delta)$-DP and has expected utility at least $\nu_{\mathcal{A}}^{f}$ over $\mathcal{M}$, then there is an algorithm $\mathcal{A}^{\prime}$ which also satisfies $(\varepsilon,\delta)$-DP and has $\nu_{\mathcal{A}^{\prime}}^{f}\geq\nu_{\mathcal{A}}^{f}$. Moreover $\mathcal{A}^{\prime}$ satisfies the following symmetry property: $\text{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]=\text{E}[X_{G,ab}^{\mathcal{A}^{\prime}}]$ for all $uv,ab\not\in E(G)$.
Proof.
Consider the following algorithm $\mathcal{A}^{\prime}$. We first choose a permutation $\pi$ of $V$ uniformly at random from the space $S_{V}$ of all permutations of $V$. Let $\pi(G)$ denote the graph isomorphic to $G$ obtained by applying this permutation, i.e., by creating an edge $\{\pi(u),\pi(v)\}$ for each edge $\{u,v\}\in E(G)$. We run $\mathcal{A}$ on $\pi(G)$, and let $\pi(H)$ denote the resulting fractional $b$-matching. We then “undo” $\pi$ to get the fractional $b$-matching $X_{G,uv}^{\mathcal{A}^{\prime}}=X_{\pi(G),\pi(u)\pi(v)}^{\mathcal{A}}$. We return $X_{G}^{\mathcal{A}^{\prime}}$.
Now we claim that $\nu_{\mathcal{A}^{\prime}}\geq\nu_{\mathcal{A}}$. Since $\pi$ was chosen uniformly at random, we know that $\pi(G)$ is distributed uniformly among $\mathcal{M}$. Hence for any $G\in\mathcal{M}$ we have that
$\text{E}[\nu_{\mathcal{A}^{\prime}}^{f}(G)]$ $=\frac{1}{n!}\sum_{\pi\in S_{V}}\text{E}[\nu_{\mathcal{A}}^{f}(\pi(G))]\geq\frac{1}{n!}\sum_{\pi\in S_{V}}\nu_{\mathcal{A}}^{f}=\nu_{\mathcal{A}}^{f}\,.$
Since this is true for all $G\in\mathcal{M}$, we have that $\nu_{\mathcal{A}^{\prime}}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}^{\prime}}(G)]\geq\nu_{\mathcal{A}}$, as claimed.
$\text{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]$ $=\frac{1}{n!}\sum_{\pi\in S_{V}}\text{E}[X_{\pi(G),\pi(u)\pi(v)}^{\mathcal{A}}]$
$=\frac{1}{n!}\sum_{G^{\prime}\in\mathcal{M}}\sum_{\pi\in\Pi(G,G^{\prime})}\text{E}[X_{G^{\prime},\pi(u)\pi(v)}^{\mathcal{A}}]$
$=\frac{1}{n!}\sum_{G^{\prime}\in\mathcal{M}}\sum_{ab\not\in E(G^{\prime})}\text{E}[X_{G^{\prime},ab}^{\mathcal{A}}].$
Hence $\text{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]$ is actually independent of $u,v$, and thus for all $uv,ab\not\in E(G)$, we have that $\text{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]=\text{E}[X_{G,ab}^{\mathcal{A}^{\prime}}]$ as claimed. ∎
Lemma B.1 implies that if we can prove an upper bound on the expected utility for DP algorithms that obey the symmetry property, then the same upper bound applies to all DP algorithms. We will now prove such a bound.
Lemma B.2.
Let $\mathcal{A}$ be an algorithm which is $(\varepsilon,\delta)$-DP, always outputs a fractional $b$-matching, and satisfies the symmetry property of Lemma B.1. Then
$\nu_{\mathcal{A}}^{f}\leq e^{2\varepsilon}b+\delta n.$
Proof.
Fix some graph $G\in\mathcal{M}$, and let $uv,ab\in E(G)$. Let $G^{\prime}$ be the graph obtained by removing $uv$ and $ab$, and adding $ub$ and $av$. Note that $G$ and $G^{\prime}$ are distance $2$ away from each other.
For all $y\neq u,b$, note that $uy\notin E(G^{\prime})$. Thus it must be the case that $\text{E}[X_{G,uy}^{\mathcal{A}}]\leq b/(n-2)$, since otherwise the symmetry property of Lemma B.1 and linearity of expectations implies that the expected (fractional) degree of $u$ in the output of $\mathcal{A}(G)$ is at least $\sum_{y\neq u,b}\text{E}[X_{G,uy}^{\mathcal{A}}]>(n-2)\frac{b}{n-2}=b$, contradicting our assumption that $\mathcal{A}$ always outputs a fractional $b$-matching. Thus $\text{E}[X_{G^{\prime},uv}^{\mathcal{A}}]\leq b/(n-2)$. Since $G$ and $G^{\prime}$ are at distance $2$ from each other, this implies that
$\text{E}[X_{G,uv}^{\mathcal{A}}]$ $=\int_{0}^{1}x\cdot\Pr[X_{G,uv}^{\mathcal{A}}=x]\mathop{}\!\mathrm{d}x$
$\leq\int_{0}^{1}x\left(e^{2\varepsilon}\cdot\Pr[X_{G^{\prime},uv}^{\mathcal{A}}=x]+2\delta\right)\mathop{}\!\mathrm{d}x$
$=e^{2\varepsilon}\cdot\text{E}[X_{G^{\prime},uv}^{\mathcal{A}}]+\delta$
$\leq e^{2\varepsilon}\frac{b}{n-2}+\delta\,.$
Since this argument applies to all $uv\in E(G)$, we have that
$\text{E}[\nu_{\mathcal{A}}^{f}(G)]$ $=\sum_{uv\in E(G)}\text{E}[X_{G,uv}^{\mathcal{A}}]$
$\leq\sum_{uv\in E(G)}\left(e^{2\varepsilon}\frac{b}{n-2}+\delta\right)$
$\leq e^{2\varepsilon}\frac{bn}{2(n-2)}+\delta\frac{n}{2}.$
$\leq e^{2\varepsilon}b+\delta n.$
Since $\nu_{\mathcal{A}}^{f}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}}^{f}(G)]$, this implies that $\nu_{\mathcal{A}}^{f}\leq e^{2\varepsilon}b+\delta n$, as claimed. ∎
We can now use these two lemmas to prove Theorem 1.2.
Proof of Theorem 1.2.
The combination of Lemma B.1 and Lemma B.2 imply that any $(\varepsilon,\delta)$-DP algorithm $\mathcal{A}$ which always outputs a fractional $b$-matching has $\nu_{\mathcal{A}}^{f}\leq e^{2\varepsilon}b+\delta n$. We also know that $\nu^{f}(G)=n/2$ for all $G\in\mathcal{M}$. Thus by the definition of an $(\alpha,\beta)$-approximation, it must be the case that $\alpha(e^{2\varepsilon}b+\delta n)+\beta\geq n/2$. ∎